# Data-informativity conditions for structured linear systems with implications for dynamic networks

### v3

@@ 8.1 Single module data-informativity results
Let ${\color[rgb]{0,0,0}\{}w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}{\color[rgb]{0,0,0}\}}$ be defined as a minimum disconnecting set from $w_{i}$ to the other components in ${\color[rgb]{0,0,0}\{}w_{\!{\scriptscriptstyle\mathcal{D}}}^{[j]}{\color[rgb]{0,0,0}\}}$, under the constraint that $i\notin\mathcal{D}^{c}$.
@@
The external signals $\chi$ that are excluded from serving as excitation sources for estimating $G_{ji}^{0}(q)$, now also include all $u$-signals that are input to a parametrized module towards the target output $w_{j}$, i.e. the inputs that correspond to parametrized terms in $\bar{T}_{(}q,\theta)]_{j*}$, the $j$-th row of $\bar{T}(q,\theta)$ in (22). Note however that, different from the term $\xi_{\!{\scriptscriptstyle\mathcal{Y}}}^{[j]}$, the signals in $u^{[j]}$ do not require excitation, as they are not part of the vector signal $\eta^{[j]}$.
@@
Then the data informativity condition (a) in Theorem 8 is generically satisfied if there exist $dim(\mathcal{D}^{c})+1$ vertex disjoint paths from ${\color[rgb]{0,0,0}\{}e^{\perp\!{\scriptscriptstyle\mathcal{Y}}},u^{\perp j}{\color[rgb]{0,0,0}\}}$ to ${\color[rgb]{0,0,0}\{}w_{i\cup\!{\scriptscriptstyle\mathcal{D}}^{c}}{\color[rgb]{0,0,0}\}}$, that do not pass through ${\color[rgb]{0,0,0}\{}w_{\!{\scriptscriptstyle\mathcal{T}}}{\color[rgb]{0,0,0}\}}$, and the signals in ${\color[rgb]{0,0,0}\{}e^{\perp\!{\scriptscriptstyle\mathcal{Y}}},u^{\perp j}{\color[rgb]{0,0,0}\}}$ are mutually independent and persistently exciting.
@@
The result shows that for appropriately exciting $w_{i\cup\!{\scriptscriptstyle\mathcal{D}}^{c}}$, we can use all external signals that in the predictor model do not have a parametrized link to the output. External signals that do have such a parametrized link are required for exciting the concerned entries in $\bar{H}(q,\theta)$ and $\bar{T}(q,\theta)$ respectively, and therefore are not available for exciting the node inputs for estimating $G_{ji}^{0}(q)$. From the remaining set of external signals, only those external signals can be used that reach the predictor input signals through the set of nodes ${\color[rgb]{0,0,0}\{}w_{i\cup\!{\scriptscriptstyle\mathcal{D}}^{c}}{\color[rgb]{0,0,0}\}}$, without passing through ${\color[rgb]{0,0,0}\{}w_{\!{\scriptscriptstyle\mathcal{T}}}{\color[rgb]{0,0,0}\}}$.
@@ 9.2 Six-node example
Given that the target module is $G_{12}$ we start composing the predictor model with input $w_{2}$ and output $w_{1}$. Because of the parallel path connection between $w_{2}$ and $w_{1}$ through $w_{3}$, we add $w_{3}$ as predictor input. And in order to handle the confounding variable that is present through the correlated disturbances $v_{1},v_{2}$, we add $\color[rgb]{0,0,0}w_{2}$} to the predictor output. This leads to a predictor model $(w_{2},w_{3})\rightarrow(w_{1},w_{2})$, that satisfies the structural conditions for consistently estimating the target module. With this choice, $w_{i}=w_{2}$ and $w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}=w_{3}$.
@@ 10 Conclusions
$\begin{bmatrix}\Delta G_{ji}(q,\theta)&\Delta_{j}^{a}(q,\theta)\!+\!\Delta_{j}^{b}(q,\theta)T_{s}(q)&\Delta_{j}^{b}(q,\theta)R_{s}(q)\end{bmatrix}\cdot$
$\cdot\begin{bmatrix}u_{i}\\ u_{\!{\scriptscriptstyle\mathcal{D}}^{c}}\\ {\color[rgb]{0,0,0}x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}\end{bmatrix}$.
Denote $\nu:=\begin{bmatrix}u_{i}\\ u_{\!{\scriptscriptstyle\mathcal{D}}^{c}}\end{bmatrix}$. Then we can decompose $\nu$ according to $\nu={\color[rgb]{0,0,0}\nu_{x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}}+{\color[rgb]{0,0,0}\nu_{\perp x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}}$ and using the fact that $\Delta_{j}^{b}(q,\theta)$ is parametrized independently from $\Delta G_{ji}$ and $\Delta_{j}^{a}$, we can rewrite $\Delta\bar{M}_{j}^{u}(q,\theta)\kappa_{u}^{[j]}$ as
$=\begin{bmatrix}\Delta G_{ji}(q,\theta)&\Delta_{j}^{c}(q,\theta)&|&\Delta_{j}^{b}(q,\theta)R_{s}(q)\end{bmatrix}\cdot$
$\cdot\begin{bmatrix}{\color[rgb]{0,0,0}\nu_{x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}}+{\color[rgb]{0,0,0}\nu_{\perp x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}}\\ {\color[rgb]{0,0,0}x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}\end{bmatrix}$
where $\Delta_{j}^{c}(q,\theta)$ replaces the term $\Delta_{j}^{a}(q,\theta)+\Delta_{j}^{b}(q,\theta)T_{s}(q)$. The expression for $\Delta\bar{M}_{j}^{u}(q,\theta)\kappa_{u}^{[j]}$ is now written as a summation of two terms that are mutually uncorrelated: one driven by $\color[rgb]{0,0,0}x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}$} and $\color[rgb]{0,0,0}\nu_{x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}$} and one driven by its orthogonal complement. According to (33), for evaluating $\Delta\bar{M}_{j}(q,\theta)\kappa^{[j]}(t)$ a third term is added, given by $\Delta\bar{M}_{j}^{e}(q,\theta)\kappa_{e}^{[j]}$, leading to
$\Delta\bar{M}_{j}(q,\theta)\kappa^{[j]}(t)$ $=$ $\begin{bmatrix}\Delta G_{ji}(q,\theta)&\Delta_{j}^{c}(q,\theta)&\!|\!&\Delta_{j}^{b}(q,\theta)R_{s}(q)\end{bmatrix}\cdot$
$\left(\begin{bmatrix}{\color[rgb]{0,0,0}v_{x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}}\\ {\color[rgb]{0,0,0}x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}\end{bmatrix}+\begin{bmatrix}{\color[rgb]{0,0,0}v_{\perp x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}}\\ 0\end{bmatrix}\right)+\Delta\bar{M}_{j}^{e}(q,\theta)\kappa_{e}^{[j]}.$
Since both terms $\color[rgb]{0,0,0}v_{\perp x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}$} and $\kappa_{e}^{[j]}$ are orthogonal to $x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}$, it follows now that $\bar{\mathbb{E}}\|\Delta\bar{M}_{j}(q,\theta)\kappa^{[j]}(t)\|^{2}=0$ implies $\begin{bmatrix}\Delta G_{ji}(q,\theta)&\Delta_{j}^{c}(q,\theta)&\Delta\bar{M}_{j}^{e}(q,\theta)\end{bmatrix}=0$ under the condition that $\Phi_{\mu^{[j]}}(\omega)\succ 0$ for almost all $\omega$ with $\mu^{[j]}$ given by
$\mu^{[j]}=\begin{bmatrix}{\color[rgb]{0,0,0}\nu_{\perp x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}}\\ \kappa_{e}^{[j]}\end{bmatrix}$,
@@
$\Delta\bar{M}_{j}(q,\theta)\kappa^{[j]}(t)$ $=$ $\Delta\bar{M}_{j}^{w}(q,\theta)w_{\!{\scriptscriptstyle\mathcal{D}}}^{[j]}(t)+\Delta\bar{M}_{j}^{\xi}(q,\theta)\xi_{\!{\scriptscriptstyle\mathcal{Y}}}^{[j]}(t)+$ (36)
@@
As a consequence of the definition of $w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}$ and the resulting $w_{\!{\scriptscriptstyle\mathcal{T}}}$, following Definition 6, the term $\Delta{\color[rgb]{0,0,0}\bar{M}_{j}^{w}}(q,\theta)w_{\!{\scriptscriptstyle\mathcal{D}}}^{[j]}$ can be written as
$\Delta{\color[rgb]{0,0,0}\bar{M}_{j}^{w}}(q,\theta)w_{\!{\scriptscriptstyle\mathcal{D}}}^{[j]}=\begin{bmatrix}\Delta\bar{G}_{ji}(q,\theta)&\Delta_{j}^{a}(q,\theta)&\Delta_{j}^{b}(q,\theta)\end{bmatrix}\begin{bmatrix}w_{i}\\ w_{\bar{\!}{\scriptscriptstyle\mathcal{D}}^{c}}\\ w_{\!{\scriptscriptstyle\mathcal{T}}}\end{bmatrix}$ (37)
with $w_{\bar{\!}{\scriptscriptstyle\mathcal{D}}^{c}}=w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}\cap w_{\!{\scriptscriptstyle\mathcal{D}}}^{[j]}$. By complementing $w_{\bar{\!}{\scriptscriptstyle\mathcal{D}}^{c}}$ with the remaining components of $w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}$ and adding 0 entries to the accompanying $\Delta^{a}_{j}(q,\theta)$, this can be rewritten as
$\Delta{\color[rgb]{0,0,0}\bar{M}_{j}^{w}}(q,\theta)w_{\!{\scriptscriptstyle\mathcal{D}}}^{[j]}=\begin{bmatrix}\Delta\bar{G}_{ji}(q,\theta)&\tilde{\Delta}_{j}^{a}(q,\theta)&\Delta_{j}^{b}(q,\theta)\end{bmatrix}\begin{bmatrix}w_{i}\\ w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}\\ w_{\!{\scriptscriptstyle\mathcal{T}}}\end{bmatrix}$. (38)
Since according to (27), $w_{\!{\scriptscriptstyle\mathcal{T}}}$ can be written as a function of $w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}$ and $\color[rgb]{0,0,0}x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}$}, substituting this in the above expression provides
@@
$\cdot\begin{bmatrix}w_{i}\\ w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}\\ {\color[rgb]{0,0,0}x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}\end{bmatrix}$.
Denote $\nu:=\begin{bmatrix}w_{i}\\ w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}\end{bmatrix}$ and $\chi:=\{{\color[rgb]{0,0,0}x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}\cup u^{[j]}\}$ . Then we can decompose $\nu$ according to $\nu=\nu_{\chi}+\nu_{\perp\chi}$ and using the fact that $\Delta_{j}^{b}(q,\theta)$ is parametrized independently from $\Delta G_{ji}$ and $\tilde{\Delta}_{j}^{a}$, we can rewrite $\Delta\bar{M}_{j}^{w}(q,\theta)w_{\!{\scriptscriptstyle\mathcal{D}}}^{[j]}$ as
$=\begin{bmatrix}\Delta G_{ji}(q,\theta)&\Delta_{j}^{c}(q,\theta)&|&\Delta_{j}^{b}(q,\theta)R_{s}(q)\end{bmatrix}\cdot$
$\cdot\begin{bmatrix}\nu_{\chi}+\nu_{\perp\chi}\\ {\color[rgb]{0,0,0}x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}\end{bmatrix}$
The expression for $\Delta\bar{M}_{j}^{u}(q,\theta)w_{\!{\scriptscriptstyle\mathcal{D}}}^{[j]}$ is now written as a summation of two terms that are mutually uncorrelated: one driven by $\color[rgb]{0,0,0}x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}$} and $\nu_{\chi}$ and one driven by a term that is orthogonal to this. For evaluating $\Delta\bar{M}_{j}(q,\theta)\kappa^{[j]}(t)$ two additional terms are added, given by $\Delta\bar{M}_{j}^{\xi}(q,\theta)\xi_{\!{\scriptscriptstyle\mathcal{Y}}}^{[j]}$ and $\Delta\bar{M}_{j}^{u}(q,\theta)u^{[j]}$.

### v4

@@ 8.1 Single module data-informativity results
Let $\{w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}\}$ be defined as a disconnecting set from $w_{i}$ to the other components in $\{w_{\!{\scriptscriptstyle\mathcal{D}}}^{[j]}\}$, under the constraint that $i\notin\mathcal{D}^{c}$.
@@
The external signals $\chi$ that are excluded from serving as excitation sources for estimating $G_{ji}^{0}(q)$, now also include all $u$-signals that are input to a parametrized module towards the target output $w_{j}$, i.e. the inputs that correspond to parametrized terms in $\bar{T}{\color[rgb]{0,0,0}(}q,\theta)]_{j*}$, the $j$-th row of $\bar{T}(q,\theta)$ in (22). Note however that, different from the term $\xi_{\!{\scriptscriptstyle\mathcal{Y}}}^{[j]}$, the signals in $u^{[j]}$ do not require excitation, as they are not part of the vector signal $\eta^{[j]}$.
@@
Then the data informativity condition (a) in Theorem 8 is generically satisfied if there exist $dim(\mathcal{D}^{c})+1$ vertex disjoint paths from $e^{\perp\!{\scriptscriptstyle\mathcal{Y}}},u^{\perp j}\}{\color[rgb]{0,0,0}\backslash\{x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}\}}$ to $\{w_{i\cup\!{\scriptscriptstyle\mathcal{D}}^{c}}\}$, and the signals in $\{e^{\perp\!{\scriptscriptstyle\mathcal{Y}}},u^{\perp j}\}$ are mutually independent and persistently exciting.
@@
The result shows that for appropriately exciting $w_{i\cup\!{\scriptscriptstyle\mathcal{D}}^{c}}$, we can use all external signals that in the predictor model do not have a parametrized link to the output. External signals that do have such a parametrized link are required for exciting the concerned entries in $\bar{H}(q,\theta)$ and $\bar{T}(q,\theta)$ respectively, and therefore are not available for exciting the node inputs for estimating $G_{ji}^{0}(q)$. From the remaining set of external signals, only those external signals can be used that reach the predictor input signals through the set of nodes $\{w_{i\cup\!{\scriptscriptstyle\mathcal{D}}^{c}}\}$, without passing through $\{w_{\!{\scriptscriptstyle\mathcal{T}}}\}$.
@@ 9.2 Six-node example
Given that the target module is $G_{12}$ we start composing the predictor model with input $w_{2}$ and output $w_{1}$. Because of the parallel path connection between $w_{2}$ and $w_{1}$ through $w_{3}$, we add $w_{3}$ as predictor input. And in order to handle the confounding variable that is present through the correlated disturbances $v_{1},v_{2}$, we add $w_{2}$ to the predictor output. This leads to a predictor model $(w_{2},w_{3})\rightarrow(w_{1},w_{2})$, that satisfies the structural conditions for consistently estimating the target module. With this choice, $w_{i}=w_{2}$ and $w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}=w_{3}$.
@@ 10 Conclusions
Acknowledegments
The authors gratefully acknowledge discussions with Xavier Bombois that contributed to the results of this paper.
@@
$\begin{bmatrix}\Delta G_{ji}(q,\theta)&\Delta_{j}^{a}(q,\theta)\!+\!\Delta_{j}^{b}(q,\theta)T_{s}(q)&\Delta_{j}^{b}(q,\theta)R_{s}(q)\end{bmatrix}$
$\cdot\begin{bmatrix}u_{i}\\ u_{\!{\scriptscriptstyle\mathcal{D}}^{c}}\\ x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}\end{bmatrix}$.
Denote $\nu:=\begin{bmatrix}u_{i}\\ u_{\!{\scriptscriptstyle\mathcal{D}}^{c}}\end{bmatrix}$. Then we can decompose $\nu$ according to $\nu=\nu_{x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}+\nu_{\perp x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}$ and using the fact that $\Delta_{j}^{b}(q,\theta)$ is parametrized independently from $\Delta G_{ji}$ and $\Delta_{j}^{a}$, we can rewrite $\Delta\bar{M}_{j}^{u}(q,\theta)\kappa_{u}^{[j]}$ as
$=\begin{bmatrix}\Delta G_{ji}(q,\theta)&\Delta_{j}^{c}(q,\theta)&|&\Delta_{j}^{b}(q,\theta)R_{s}(q)\end{bmatrix}\begin{bmatrix}\nu_{x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}\!+\!\nu_{\perp x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}\\ x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}\end{bmatrix}$
where $\Delta_{j}^{c}(q,\theta)$ replaces the term $\Delta_{j}^{a}(q,\theta)+\Delta_{j}^{b}(q,\theta)T_{s}(q)$. The expression for $\Delta\bar{M}_{j}^{u}(q,\theta)\kappa_{u}^{[j]}$ is now written as a summation of two terms that are mutually uncorrelated: one driven by $x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}$ and $\nu_{x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}$ and one driven by its orthogonal complement. According to (33), for evaluating $\Delta\bar{M}_{j}(q,\theta)\kappa^{[j]}(t)$ a third term is added, given by $\Delta\bar{M}_{j}^{e}(q,\theta)\kappa_{e}^{[j]}$, leading to
$\Delta\bar{M}_{j}(q,\theta)\kappa^{[j]}(t)\!=\!\begin{bmatrix}\Delta G_{ji}(q,\theta)&\Delta_{j}^{c}(q,\theta)&\!|\!\!&\Delta_{j}^{b}(q,\theta)R_{s}(q)\end{bmatrix}$
$\hskip 56.9055pt\cdot\left(\begin{bmatrix}v_{x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}\\ x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}\end{bmatrix}+\begin{bmatrix}v_{\perp x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}\\ 0\end{bmatrix}\right)+\Delta\bar{M}_{j}^{e}(q,\theta)\kappa_{e}^{[j]}.$
Since both terms $v_{\perp x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}$ and $\kappa_{e}^{[j]}$ are orthogonal to $x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}$, it follows now that $\bar{\mathbb{E}}\|\Delta\bar{M}_{j}(q,\theta)\kappa^{[j]}(t)\|^{2}=0$ implies $\begin{bmatrix}\Delta G_{ji}(q,\theta)&\Delta_{j}^{c}(q,\theta)&\Delta\bar{M}_{j}^{e}(q,\theta)\end{bmatrix}=0$ under the condition that $\Phi_{\mu^{[j]}}(\omega)\succ 0$ for almost all $\omega$ with $\mu^{[j]}$ given by
$\mu^{[j]}=\begin{bmatrix}\nu_{\perp x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}}\\ \kappa_{e}^{[j]}\end{bmatrix}$,
@@
$\Delta\bar{M}_{j}(q,\theta)\kappa^{[j]}(t)$ $=$ $\Delta\bar{M}_{j}^{w}(q,\theta)w_{\!{\scriptscriptstyle\mathcal{D}}}^{[j]}(t)\!+\!\Delta\bar{M}_{j}^{\xi}(q,\theta)\xi_{\!{\scriptscriptstyle\mathcal{Y}}}^{[j]}(t)$ (36)
@@
As a consequence of the definition of $w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}$ and the resulting $w_{\!{\scriptscriptstyle\mathcal{T}}}$, following Definition 6, the term $\Delta\bar{M}_{j}^{w}(q,\theta)w_{\!{\scriptscriptstyle\mathcal{D}}}^{[j]}$ can be written as
$\Delta\bar{M}_{j}^{w}(q,\theta)w_{\!{\scriptscriptstyle\mathcal{D}}}^{[j]}=\begin{bmatrix}\Delta\bar{G}_{ji}(q,\theta)&\Delta_{j}^{a}(q,\theta)&\Delta_{j}^{b}(q,\theta)\end{bmatrix}\begin{bmatrix}w_{i}\\ w_{\bar{\!}{\scriptscriptstyle\mathcal{D}}^{c}}\\ w_{\!{\scriptscriptstyle\mathcal{T}}}\end{bmatrix}$
with $\{w_{\bar{\!}{\scriptscriptstyle\mathcal{D}}^{c}}\}=\{w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}\}\cap\{w_{\!{\scriptscriptstyle\mathcal{D}}}^{[j]}\}$. By complementing $w_{\bar{\!}{\scriptscriptstyle\mathcal{D}}^{c}}$ with the remaining components of $w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}$ and adding 0 entries to the accompanying $\Delta^{a}_{j}(q,\theta)$, this can be rewritten as
$\Delta\bar{M}_{j}^{w}(q,\theta)w_{\!{\scriptscriptstyle\mathcal{D}}}^{[j]}=\begin{bmatrix}\Delta\bar{G}_{ji}(q,\theta)&\tilde{\Delta}_{j}^{a}(q,\theta)&\Delta_{j}^{b}(q,\theta)\end{bmatrix}\!\begin{bmatrix}w_{i}\\ w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}\\ w_{\!{\scriptscriptstyle\mathcal{T}}}\end{bmatrix}$.
Since according to (27), $w_{\!{\scriptscriptstyle\mathcal{T}}}$ can be written as a function of $w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}$ and $x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}$, substituting this in the above expression provides
@@
$\begin{bmatrix}w_{i}\\ w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}\\ x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}\end{bmatrix}$.
Denote $\nu:=\begin{bmatrix}w_{i}\\ w_{\!{\scriptscriptstyle\mathcal{D}}^{c}}\end{bmatrix}$ and $\chi:=\{x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}\cup u^{[j]}\}$ . Then we can decompose $\nu$ according to $\nu=\nu_{\chi}+\nu_{\perp\chi}$ and using the fact that $\Delta_{j}^{b}(q,\theta)$ is parametrized independently from $\Delta G_{ji}$ and $\tilde{\Delta}_{j}^{a}$, we can rewrite $\Delta\bar{M}_{j}^{w}(q,\theta)w_{\!{\scriptscriptstyle\mathcal{D}}}^{[j]}$ as
$\begin{bmatrix}\Delta G_{ji}(q,\theta)&\Delta_{j}^{c}(q,\theta)&|&\Delta_{j}^{b}(q,\theta)R_{s}(q)\end{bmatrix}\!\cdot\!\begin{bmatrix}\nu_{\chi}+\nu_{\perp\chi}\\ x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}\end{bmatrix}$
The expression for $\Delta\bar{M}_{j}^{u}(q,\theta)w_{\!{\scriptscriptstyle\mathcal{D}}}^{[j]}$ is now written as a summation of two terms that are mutually uncorrelated: one driven by $x_{\!{\scriptscriptstyle\mathcal{T}}^{*}}$ and $\nu_{\chi}$ and one driven by a term that is orthogonal to this. For evaluating $\Delta\bar{M}_{j}(q,\theta)\kappa^{[j]}(t)$ two additional terms are added, given by $\Delta\bar{M}_{j}^{\xi}(q,\theta)\xi_{\!{\scriptscriptstyle\mathcal{Y}}}^{[j]}$ and $\Delta\bar{M}_{j}^{u}(q,\theta)u^{[j]}$.